If the unmodulated current in the antenna load of an AM transmitter is 10 A, determine the increase in this current resulting from 80% modulation.
Answer #1

Answer: The unmodulated current is 10A and modulation index is .8. so apply the formula of current of increase in current as I=Ic\sqrt{}1+.82/2

Detailed solution is attached here

  Ita Icllt u M=.8 2 = 10 /1+0.83 Elo (1+032) = 10 (1:14) It = 1114 Scanned with CamScanner